OpenStudy (anonymous):

Solve 4x + 6 = 38 (Points : 4) 144 11 8 3.5

OpenStudy (anonymous):

Well, the first thing you need to do is isolate the variable on one side. To do this, subtract six from both sides. 4x + 6 = 38 - 6 -6 4x = 32 Now, you need to isolate just x. In order to do this, you need to divide each side by four. 4x/4 = 32/4 What do you get?